OpenStudy (anonymous):

The center of a circle with the equation (x - 6)2 + (y + 2)2 = 49 is ( 6, -2). A. True B. False

OpenStudy (anonymous):

i got this

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

True

OpenStudy (anonymous):

how do u know?

OpenStudy (anonymous):

yea i don't got this XD

OpenStudy (anonymous):

http://prntscr.com/3co9wk

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

lol, you're welcome ^.^ I actually know how to do this cx

OpenStudy (anonymous):

A low-watt radio station can be heard only within a certain distance from the station. On the graph below, the circular region represents that part of the city where the station can be heard, and the center of the circle represents the location of the station. Which equation represents the boundary for the region where the station can be heard? A. (x – 4)2 + (y – 5)2 = 50 B. (x + 5)2 + (y – 5)2 = 59 C. (x + 5)2 + (y + 4)2 = 25 D. (x + 5)2 + (y – 5)2 = 25

OpenStudy (anonymous):

@Vallarylynn16

OpenStudy (anonymous):

@Vallarylynn16 hey, hey, better not be talking about me XD

OpenStudy (anonymous):

lol no, I'm not.. give me a min.. I have to graph them all cx

OpenStudy (anonymous):

Either B or C

OpenStudy (anonymous):

wich is the best answer

OpenStudy (anonymous):

the both are within range... B is a bit closer.. so try that one..?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

it was c

OpenStudy (anonymous):

/.\ sorry lol

OpenStudy (anonymous):

but thanks

OpenStudy (anonymous):

you good now..? Can I leave and not get my notifications blown up..? cx

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

For the coordinate plane, which equation describes the locus of points equidistant from (–1, 3) and (3, –1)? A. y = -x B. y + 1 = x – 3 C. y – 3 = –(x + 1) D. y = x

OpenStudy (anonymous):

@Vallarylynn16

OpenStudy (anonymous):

what is the locus of points..?

OpenStudy (anonymous):

anyway.. the only option that runs through those points is y-3=-(x+1)

OpenStudy (anonymous):

http://prntscr.com/3cz8h5

OpenStudy (anonymous):

so the answer is c

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I'm not here to give you the answer, I'm here to help you find it lol

OpenStudy (anonymous):

..... really

OpenStudy (anonymous):

your gonna really pull that on me

OpenStudy (anonymous):

lol yes, yes I am, I believe the answer is pretty obvious w/o me telling you

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

YES OR NO

OpenStudy (anonymous):

OMG JUST CLICK ON IT cx

OpenStudy (anonymous):

i did! lawlz

OpenStudy (anonymous):

A circle has radius 6 cm. In the plane of the circle, what best describes the locus of points that are 2 cm from the circle? A. a circle of radius 2 cm B. a circle of radius 4 cm and a circle of radius 8 cm C. a circle of radius 8 cm D. a circle of radius 2 cm and a circle of radius 8 cm

OpenStudy (anonymous):

@Vallarylynn16

OpenStudy (anonymous):

I have no idea ...

OpenStudy (anonymous):

do you know what the 'locus of points' is..? that might help lol

OpenStudy (anonymous):

i got it.. nvrm